# ModelNet40-E: An Uncertainty-Aware Benchmark for Point Cloud Classification

### v1

@@ Introduction
PointNet 89.30 86.30 86.18 84.64
DGCNN 91.00 89.67 89.34 89.02
PTv3 88.82 88.74 88.90 88.05
Expected Calibration Error (ECE)
Figure 4: Accuracy (left), expected calibration error (center), and Pearson correlation between predicted and measurement uncertainty (right) for PTv3 across training epochs under varying noise levels.
We evaluate the Expected Calibration Error (ECE) of all models under varying noise levels, following the procedure described in Section 2.
The results, summarized in Table 3, show that calibration error increases consistently with noise severity. This indicates that higher measurement uncertainty not only degrades classification accuracy but also results in less reliable confidence estimates. Among the evaluated architectures, PTv3 consistently achieves the lowest ECE across all noise configurations, suggesting superior calibration performance compared to both PointNet and DGCNN.
@@
PointNet 0.0561 0.0821 0.0848 0.0994
DGCNN 0.0493 0.0522 0.0553 0.0591
PTv3 0.0388 0.0223 0.0314 0.0419
[…]
To further investigate the relationship between input noise and calibration, we stratify the test samples into quartiles based on their mean measurement uncertainty $\sigma$ and report the Expected Calibration Error. For each model, we collect the values of $\sigma$ from all noise levels and divide them into four bins of equal size, ranging from the lowest to the highest uncertainty. Results are presented in Table 4.
For all models, ECE increases monotonically from Bin 0 (lowest uncertainty) to Bin 3 (highest uncertainty), confirming that measurement noise negatively impacts model calibration. While DGCNN maintains relatively stable ECE across bins, PointNet shows a more pronounced degradation in calibration with increasing noise. PTv3 again demonstrates the lowest ECE in all quartiles, indicating not only strong overall calibration but also robustness to varying levels of measurement uncertainty.
Table 4: Stratified expected calibration error computed across quartiles of measurement uncertainty $\sigma$. Bin 0 corresponds to the lowest $\sigma$ quartile and Bin 3 to the highest. Lower is better.
Model Bin 0 Bin 1 Bin 2 Bin 3
PointNet 0.0561 0.0799 0.0850 0.1018
DGCNN 0.0499 0.0512 0.0542 0.0616
PTv3 0.0395 0.0247 0.0308 0.0437
We plot calibration curves for each subset in Figure 3. Across all models, we observe a clear degradation in calibration as $\sigma$ increases. In the lowest quartile (Q1), models are relatively well-calibrated, with accuracy closely tracking confidence. However, from Q2 onward, calibration quality deteriorates, particularly in the low-confidence bins, where models tend to be significantly overconfident. This trend becomes more pronounced in Q3 and Q4, where curves increasingly deviate from the ideal diagonal. PTv3 maintains a modest advantage in calibration across most quartiles, while PointNet shows greater overconfidence in higher-noise regimes. These results align with the overall ECE trends and reinforce the sensitivity of model calibration to input-level measurement noise.
[…]
@@
Model Light Moderate Heavy
PointNet 0.0726 0.0379 0.0437
DGCNN 0.1346 0.0739 0.0716
PTv3 0.1506 0.1697 0.1453
Correlation Between Measurement Uncertainty and Predicted Uncertainty
@@
Table 5 reports the Pearson correlation between the ground-truth measurement uncertainty $\sigma$ and the predicted uncertainty (see Section 2).
Among the evaluated models, PTv3 consistently achieves the highest correlation across all noise levels, with values close to 0.19, suggesting that its confidence estimates are strongly aligned with the true measurement uncertainty. This indicates that PTv3 not only maintains a low Expected Calibration Error but also exhibits well-calibrated uncertainty estimates that appropriately respond to increases in noise level.
DGCNN also shows moderate positive correlation, achieving values between 0.07 and 0.13. While this is substantially higher than PointNet in all cases, the correlation decreases notably under moderate and heavy noise levels, indicating some limitations in DGCNN’s uncertainty awareness in more challenging conditions.
In contrast, PointNet demonstrates considerably lower correlation, with values near 0.04 under moderate and heavy noise, indicating that its predicted confidence is largely insensitive to the actual measurement uncertainty. These findings highlight that while all models can achieve high classification accuracy, their ability to recognize when predictions should be uncertain varies significantly. The strong performance of PTv3 suggests that modern transformer-based architectures may be better suited to capture and reflect input uncertainty, whereas simpler point-based networks are more prone to overconfident predictions regardless of noise level.
Training Dynamics for PTv3
@@
We investigate the classification accuracy, calibration (ECE), and correlation between predicted uncertainty and measurement uncertainty for PTv3 across several training times and noise levels. For each noise level, we save the best model (in terms of classification accuracy) after 20, 40, 60, 80, and 100 training epochs.
Figure 4 shows our results. As expected, classification accuracy (left) improves steadily over time. However, ECE (center) tends to worsen as training progresses, especially under moderate and heavy noise. This suggests that while the model becomes more accurate, its confidence estimates become less aligned with actual correctness, indicating degraded calibration.
Most notably, the Pearson correlation between predicted and measurement uncertainty (right) also decreases with additional training epochs. This decline implies that the model becomes less sensitive to input degradation, even as it becomes more confident and accurate. Such a trend aligns with prior observations in the literature, where overfitting can lead to overconfident predictions and reduced uncertainty awareness (Guo et al. 2017).
These findings underscore the importance of evaluating not only accuracy but also calibration and uncertainty sensitivity when training deep learning models, especially for safety-critical applications. Rather than relying solely on classification accuracy as a stopping criterion, monitoring calibration and uncertainty awareness could provide more robust indicators for determining when to halt training.
Conclusion
We present ModelNet40-E, a novel uncertainty-aware benchmark designed to evaluate the robustness, calibration, and uncertainty sensitivity of point cloud classification models. We introduce LiDAR-like noise to the original ModelNet40 dataset at three distinct noise severity levels.
We evaluate three representative architectures—PointNet, DGCNN, and Point Transformer v3 (PTv3)—across classification accuracy, Expected Calibration Error (ECE), and uncertainty awareness (correlation between measurement and predicted uncertainty). Our results reveal that classification accuracy alone is insufficient to capture model reliability under corruption. For instance, while DGCNN achieves the highest accuracy across noise levels, PTv3 consistently demonstrates superior calibration and uncertainty sensitivity. Notably, PTv3 is the only model to maintain a relatively stable Pearson correlation between measurement uncertainty and predicted uncertainty as noise severity increases, reflecting a more robust and reliable uncertainty estimation. Moreover, our analysis of training dynamics shows that prolonged training can degrade calibration and uncertainty sensitivity even as accuracy continues to improve, suggesting the importance of monitoring these metrics as potential early stopping criteria.

### v2

@@ 4.3.1 Classification Accuracy
PointNet 88.86 65.11 53.69 33.79
PointNet++ 91.61 70.91 52.84 16.21
DGCNN 88.78 60.09 50.49 36.99
SimpleView 85.62 84.60 80.47 63.01
CurveNet 89.14 69.12 58.55 38.09
PointMLP 89.99 76.01 59.08 25.32
PTv3 89.47 79.42 71.11 43.84
4.3.2 Expected Calibration Error (ECE)
@@
PointNet 0.0628 0.1897 0.2671 0.4269
PointNet++ 0.0564 0.1441 0.2629 0.6417
DGCNN 0.0708 0.2664 0.3325 0.4550
SimpleView 0.0800 0.0828 0.1048 0.2392
CurveNet 0.0635 0.1666 0.2512 0.4400
PointMLP 0.0626 0.1294 0.2613 0.5835
PTv3 0.0437 0.1004 0.1488 0.3492
Table 3 reports the Expected Calibration Error across all models and noise levels. As with accuracy, calibration deteriorates as noise severity increases, indicating that higher measurement uncertainty not only reduces predictive performance but also leads to less reliable confidence estimates.
PTv3 achieves the best calibration on the clean set, even if it did not achieve the highest accuracy. Under noise, both PTv3 and SimpleView present the lowest ECE, with SimpleView performing slightly better. Notably, while both PointNet++ and PointMLP maintain a relatively low ECE under clean and light noise levels, suffer from a sharp degradation under moderate and heavy noise levels, a behavior that is consistent with their accuracy results.
Figure 1 shows the calibration curves for all models. The dashed diagonal indicates perfect calibration, while the gray shaded area indicates low-confidence predictions ($40\%$), which are underrepresented and thus less reliable. Across all noise configurations, models tend to become increasingly overconfident as noise severity rises, with curves falling below the diagonal. This effect is particularly notable under moderate and heavy noise.
Table 4: Uncertainty awareness results across all models. We report AUROC for error detection under different noise levels (None, Light, Moderate, Heavy), and Pearson correlation between ground-truth measurement uncertainty $\sigma$ and predicted uncertainty $(1-p_{\text{max}})$. Correlations are calculated after pooling samples across all noise levels, including only correctly classified samples (correct only), and all samples (all). Higher AUROC and Pearson values indicate stronger error detection and uncertainty awareness, respectively.
@@
Model AUROC (None) AUROC (Light) AUROC (Moderate) AUROC (Heavy) Pearson (correct only) Pearson (all)
PointNet 0.8971 0.8367 0.7650 0.6584 0.1787 0.1180
PointNet++ 0.9081 0.8446 0.7949 0.5825 0.1400 0.0463
DGCNN 0.9043 0.8102 0.7782 0.6574 0.1334 0.0754
SimpleView 0.8746 0.8498 0.8306 0.7230 0.1291 0.1397
CurveNet 0.8852 0.8138 0.7353 0.5885 0.1295 0.0520
PointMLP 0.8798 0.8526 0.7309 0.5178 0.2003 0.0918
PTv3 0.8980 0.8653 0.8282 0.7292 0.1853 0.2118
4.3.3 Uncertainty Awareness
@@
Table 4 reports the results for uncertainty awareness, including (i) AUROC and (ii) Pearson correlation between ground-truth measurement uncertainty $\sigma$ and predicted uncertainty.
(1) AUROC. It measures the ability of each model to distinguish between correct and incorrect predictions using predicted uncertainty. Across all architectures, AUROC decreases as noise severity increases, indicating that error detection becomes more difficult under stronger perturbations. Among the models, PointNet++ achieves the highest AUROC on clean data, while SimpleView achieves the best performance under moderate noise. PTv3 is the most consistent overall, ranking near the top across all noise levels and achieving the highest AUROC under light and heavy noise.
(2) Correlation Between Measurement Uncertainty and Predicted Uncertainty. This metric evaluates whether models reduce their confidence as measurement noise increases. To isolate genuine uncertainty awareness from the trivial confidence drop caused by misclassifications, we report correlations on correctly classified samples as our primary metric (see Section 1). For completeness, we also include results on all samples, providing additional insights into whether models remain uncertainty-aware even when making incorrect predictions.
PointMLP achieves the strongest correlation on correctly classified samples, followed by PTv3 and PointNet, indicating that these models adjust their confidence most consistently in response to increasing noise while still making correct predictions. However, PointMLP’s correlation drops substantially when all samples are included, suggesting that while it is highly sensitive to noise on correct predictions, its uncertainty estimates do not reliably capture misclassified inputs. In contrast, PTv3 achieves both strong correct-only correlation and the highest all-sample correlation, demonstrating more consistent uncertainty-awareness, which is also supported by its superior AUROC results.
SimpleView, despite achieving excellent robustness under noise for accuracy and ECE, shows a relatively weak correlation, implying that its confidence estimates are less sensitive to input noise.
4.3.4 Training Dynamics
Figure 2: Training dynamics of accuracy (top row), Expected Calibration Error (ECE, middle row), and AUROC (bottom row) for all models across training epochs under varying noise levels.
Figure 3: Training dynamics of Pearson correlation between ground-truth measurement uncertainty and predicted uncertainty across training epochs, computed on correctly classified samples (left) and on all samples (right). Correlations are calculated after pooling samples across all noise levels.
@@ 4.3.4 Training Dynamics
We investigate the training dynamics for all models by evaluating accuracy, Expected Calibration Error (ECE), AUROC, and Pearson correlation between ground-truth and predicted uncertainty across several training times and noise levels. Each model is assessed after 20, 40, 60, 80, and 100 training epochs.
Figure 2 shows our results for accuracy, ECE, and AUROC, presented by noise level. As training progresses, accuracy generally improves across models, but this improvement often comes at the cost of higher calibration error, particularly under moderate and heavy noise. This effect is particularly evident for SimpleView. AUROC trends are more stable, though some models such as PointMLP show notable fluctuations under moderate and heavy noise levels.
Figure 3 presents the dynamics of the Pearson correlations, for both correct-only and all samples. In this case, the results are derived after pooling samples across all noise levels. On correctly classified samples, most models initially achieve moderate correlation values that gradually decline with training, suggesting a loss of sensitivity to input noise, particularly for PointNet++ and DGCNN. When including all samples, correlations are typically lower, reflecting the impact of misclassifications. Interestingly, PTv3 maintains relatively strong and stable correlations in both settings, while other models such as PointMLP show strong correct-only correlations but collapse when misclassified samples are included.
These results highlight a key trade-off: longer training improves accuracy but can degrade both calibration and uncertainty awareness. Such a trend aligns with prior observations in the literature (Guo et al., 2017). Monitoring metrics beyond accuracy—particularly ECE and correlation-based uncertainty measures—provides a more complete picture of model reliability under noise.
5 Conclusion
In this work, we introduced ModelNet40-E, an uncertainty-aware benchmark for point cloud classification. Unlike existing benchmarks that evaluate models solely on accuracy, ModelNet40-E enables systematic assessment of calibration and uncertainty awareness under LiDAR-like noise. By providing ground-truth measurement uncertainty alongside corrupted point clouds, ModelNet40-E enables a clear separation between robustness, calibration, and uncertainty awareness.
Our evaluation across multiple representative architectures highlights several key findings. First, robustness to noise varies significantly: while early models such as PointNet and DGCNN suffer sharp degradation, more recent approaches like SimpleView and PTv3 maintain higher accuracy under corruption. Second, calibration errors grow substantially with noise, but transformer-based PTv3 demonstrates consistently superior calibration compared to other models. Finally, our new correlation-based metric reveals notable differences in uncertainty awareness. PTv3 achieves both strong correct-only and all-sample correlations, reflecting reliable noise sensitivity, whereas models like PointMLP exhibit strong awareness on correct predictions but fail to capture uncertainty on misclassified samples.
LLM Usage Disclosure
This paper benefited from the use of large language models (LLMs) to assist with writing, code development, and literature research. All outputs generated by LLMs were carefully reviewed and verified by the authors, and all scientific contributions were developed by the authors.